OpenStudy (anonymous):

how do i solve this? (transforming formulas algebra 1) please show me step by step it would help alot thanks 10ac-x/11=-3 solve for a

OpenStudy (anonymous):

is it all over 11 like \[\frac{10ac-x}{11}=-3\]

OpenStudy (anonymous):

or is just the \(x\) over 11 like \[10ac-\frac{x}{11}=-3\]?

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

step one: multiply both sides by \(11\) to get rid of annoying fraction you get \[10ac-x=-33\]

OpenStudy (anonymous):

step 2: add \(x\) to both sides and get \[10ac=-33+x\] or if you prefer \[10ac=x-33\]

OpenStudy (anonymous):

and since you are solving for \(a\) the last step is to divide both sides by \(10c\) to get \[a=\frac{x-33}{10c}\] then it is done

OpenStudy (anonymous):

that was step by step as requested, hope it is okay

OpenStudy (anonymous):

thank you so much, it was very helpful!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

 what about this one?